Bill is one year less than twice as old as his sister Caroline. The total of their ages is 26. How old is Bill?
Answer by JBarnum(2146)

Bill and Caroline
B=2C-1
B+C=26
2C-1+C=26
3C-1=26
3C=27
C=9
B=17